# Asymmetries and Circumstellar Interaction in the Type II SN 2024bch

### v1

@@ Abstract
We present a comprehensive multi-epoch photometric and spectroscopic study of SN 2024bch, a nearby (19.9 Mpc) Type II supernova (SN) with prominent early high ionization emission lines. Optical spectra from 2.9 days after the estimated explosion reveal narrow lines of H I, He II, C IV, and N IV that disappear by day 6. High cadence photometry from the ground and TESS show that the SN brightened quickly and reached a peak M 17.8 mag within a week of explosion, and late-time photometry suggests a 56Ni mass of 0.050 M☉. High-resolution spectra from day 8 and 43 trace the unshocked circumstellar medium (CSM) and indicate a wind velocity of 30–40 km s-1, a value consistent with a red supergiant (RSG) progenitor. Comparisons between models and the early spectra suggest a pre-SN mass-loss rate of , which is too high to be explained by quiescent mass loss from RSGs, but is consistent with some recent measurements of similar SNe. Persistent blueshifted H I and [O I] emission lines seen in the optical and NIR spectra could be produced by asymmetries in the SN ejecta, while the multi-component H may indicate continued interaction with an asymmetric CSM well into the nebular phase. SN 2024bch provides another clue to the complex environments and mass-loss histories around massive stars.
@@ 1 Introduction
Extensive mass-loss is an important property of massive stars, and in many cases it appears to ramp up to extreme levels in the years to decades before core collapse (Smith, 2014). While the final stages of a massive star’s life may be hard to observe, some of the information is recorded in the surrounding stellar envelope and circumstellar material (CSM) and can be observed during the first few hours and days of a core collapse supernova (CCSN) explosion. This ‘flash spectroscopy’ includes fleeting narrow lines of H, He, C, and N. While this phenomenon had been well documented in several lucky cases described in the literature (Niemela et al., 1985; Benetti et al., 1994; Leonard et al., 2000; Quimby et al., 2007; Shivvers et al., 2015), within the last decade the numbers of CCSN observed early enough to detect interaction with the mass-loss from the progenitor has dramatically increased thanks to fast response by dedicated transient surveys (for example: Gal-Yam et al., 2014; Smith et al., 2015; Yaron et al., 2017; Khazov et al., 2016; Bullivant et al., 2018; Hosseinzadeh et al., 2018; Tartaglia et al., 2021; Bruch et al., 2021; Terreran et al., 2022; Jacobson-Galán et al., 2022; Andrews et al., 2024). The recent and very nearby ( 7 Mpc) SNe 2023ixf and 2024ggi alone have contributed many publications analyzing their early-time luminosity evolution and exquisite early time spectroscopic datasets (Smith et al., 2023; Li et al., 2024; Zimmerman et al., 2024; Jacobson-Galán et al., 2023; Bostroem et al., 2023a; Hiramatsu et al., 2023; Teja et al., 2023; Zhang et al., 2023; Shrestha et al., 2024a; Jacobson-Galán et al., 2024b; Zhang et al., 2024; Pessi et al., 2024; Dickinson et al., 2024).
@@
The CCSNe that result from a progenitor that retains most of the hydrogen envelope prior to explosion are classified as Type II (see Arcavi, 2017; Gal-Yam, 2017; Branch & Wheeler, 2017, for detailed reviews). Some of these objects, particularly those with plateau light curves (SNe II-P), have been confirmed to come from red supergiant (RSG) progenitor stars via Hubble Space Telescope imaging (Smartt, 2015; Van Dyk, 2015). When Type II SNe are observed early enough (hours to a few days after explosion) we can often observe shock interaction between the ejecta and the CSM or from the shock breakout of the SN itself (Bruch et al., 2021). These fleeting emission lines show Lorentzian profiles that can have wings extending out to 1500 - 2000 km s-1 due to electron scattering in the ionized gas (see Smith, 2017 for a review of SNe with dense CSM). Early spectroscopy can be a powerful device for understanding this enhanced mass loss and its potential mechanisms prior to collapse by probing the density, temperature, and composition of the surrounding CSM.
@@
While eruptive mass loss in luminous blue variables (LBVs) and strong line-driven winds of Wolf-Rayet (WR) stars are well known (see Smith, 2014 for a review), for RSGs there have been significant reductions in wind mass-loss prescriptions, and significant mismatch between these expected RSG wind mass-loss rates and mass-loss rates calculated from CSM interaction in SNe. For instance, the recent RSG wind mass-loss prescription from Beasor et al. (2020), which is estimated from modern high-quality IR data, reduces RSG mass-loss rates by about an order of magnitude compared to the older 1970s and 1980s literature-based canonical prescription from de Jager et al. (1988). Some stellar evolution models invoke even higher “enhanced” mass-loss rates (Ekström et al., 2012), but even these artificially inflated rates are still not enough to explain the estimated via flash spectroscopy observations (Jacobson-Galán et al., 2024a). This discrepancy could be resolved if periods of enhanced pre-SN mass loss are considered for RSGs rather than steady, quiescent mass loss from winds.
@@
Causes for increased periods of mass loss could be from wave driven mass loss during late-stage nuclear burning (Quataert & Shiode, 2012; Fuller, 2017; Wu & Fuller, 2021), turbulent convection in the core due to some dynamical instability (Smith & Arnett, 2014), or even interaction with a companion and a SN progenitor’s inflated envelope (Smith & Arnett, 2014), as the binary fraction for massive stars is high (Sana et al., 2012; Moe & Di Stefano, 2017). Eruptive mass loss from the progenitors of normal Type II SNe have been very difficult to detect observationally (Kochanek et al., 2017; Dong et al., 2023; Shrestha et al., 2024a, for example) suggesting if they occur they are likely some combination of dusty, faint, and quick, although see Jacobson-Galán et al. (2022) for one possible exception.
@@ 2 Discovery, Distance, and Reddening
SN 2024bch is located in NGC 3206 (Figure 1) at J2000 coordinates , . It was discovered on 2024-01-29.25 (Wiggins, 2024, MJD 60338.25) and classified later that night as a possible IIn (n is for “narrow” spectral lines), resembling the well-studied IIn SN 1998S (Balcon, 2024). While there are similarities between the two spectra, the further evolution of SN 2024bch indicated that its IIn-like features faded more quickly than SN 1998S, and that the narrow lines were likely those typically seen in objects exhibiting flash ionization. A nondetection of the object was reported by the Gravitational-wave Optical Transient Observer (Steeghs et al., 2022, GOTO) on 2024-01-28.04 (MJD 60337.04) with a limiting magnitude of 18.2 in their L-GOTO filter. A day later on 2024-01-29.06 (MJD 60338.06) GOTO recorded a detection at the same magnitude of L = 18.2 0.06, which was observed a few hours before the data cited in the discovery report. Without any other observations during this time to help constrain the date of explosion we will use the midway point between the GOTO limiting magnitude and first detection, or MJD = 60337.55, to be our explosion epoch which we will use throughout this paper.
@@
| First Detection (JD) | 2460338.56 |
| Explosion Epoch (JD)aamid point of last nondetection and first detection | 2460338.05 |
@@ 4.1 Optical and UV Lightcurves
The full optical and UV lightcurves of SN 2024bch are shown in Figure 3, and the V-band lightcurve in absolute magnitude is shown in the bottom of Figure 4. Our estimated explosion epoch of MJD = 60337.55 is indicated by the dashed line in the right panel of Figure 3. We were able to observe fairly continuously from 1.5 days out to 155 days before the SN became Sun constrained. From our data it appears that the V band reaches a maximum mag at roughly 7 days post explosion, while the and -band takes much longer, roughly 15 to 17 days to reach the maximum. This could potentially be due to CSM interaction that will be discussed in detail below, but is still within somewhat normal values for some Type II SNe (González-Gaitán et al., 2015). The observations also show a peak at roughly 5 days, once again on the shorter end of the distribution, but within the normal range. We also catch a rise in the Swift data between our first and second epoch indicating maximum was reached somewhere near day 3. Unfortunately, the cadence is not very high for our Swift observations and our first epoch was at 2.3 days after explosion so we cannot constrain the UV behavior more than this.
As can be seen in Figure 3, the plateau length persists only until 70 days post-explosion, which is a bit on the shorter side for Type II-P supernovae (Valenti et al., 2016). The fall from the plateau lasts about 15-20 days (the observational cadence was quite low during this time), and when the radioactive tail phase begins mag, or about 3 magnitudes fainter than at peak (bottom panel Figure 4). Overall the shape and brightness of the V-band lightcurve is quite similar to the recent and well-studied Type II SN 2023ixf, both in plateau duration and shape and in the post-plateau brightness (Hsu et al., 2024). The lightcurves of SN 2024ggi and SN 2022jox, other recent CCSNe showing flash features, also have very similar properties at least until about 50 days when the lightcurve of SN 2024ggi shows a much more pronounced and prolonged plateau. SN 2017gmr and SN 2017eaw, and the second half of the plateau phase for SN 2024ggi, show the more classical plateau shape and length. Interestingly, both SN 2017gmr and SN 2017eaw were discovered quite early and have spectroscopy within the first few days and do not show flash features other than possibly fleeting narrow H and broad He bumps (Andrews et al., 2019; Rui et al., 2019), and SN 2024ggi only showed flash features until day 3-4 (Shrestha et al., 2024a; Zhang et al., 2024; Jacobson-Galán et al., 2024b).
@@ 4.2 Color Evolution
In the top panel of Figure 4 we show the color evolution of SN 2024bch compared with other well studied Type II SNe, with a zoom in over the first few days shown in Figure 5. As seen from the figures, we are able to catch a blueward color evolution over the first three days, reaching mag, and then a slow evolution to the red until reaching a value of = 1.0 mag around day 75, when the fall from plateau begins. From this point onward there is a quick drop to = 0.9 mag and the color slowly trends blueward with a = 0.8 mag around day 150 when the SN becomes Sun constrained.
@@ 4.3 Bolometric Lightcurve and 56Ni Mass
The late-time bolometric luminosity appears to decline at a rate of 0.014 mag d-1, a value higher than the canonical rate of fully-trapped 56Co decay of 0.0098 mag d-1. This is shown as the dashed line in the bottom of Figure 6 where we show the decay of a mass of 56Ni = 0.055 M☉ (a value we derive below) for comparison. The slight deviation and quicker decline of SN 2024bch may indicate that the gamma rays were not fully trapped, that there is potentially dust formation occurring, or that there is CSM interaction which is contributing to the overall luminosity and not only increasing the luminosity but fading at a higher rate than radioactive decay alone. While we will explore these in more detail later, average values for 56Ni masses in normal Type II SNe have been found to be 0.04 M☉ (Valenti et al., 2016; Müller et al., 2017; Anderson, 2019; Rodríguez et al., 2021). The slightly larger estimated 56Ni mass combined with the shape of the H emission (described in section 5.3) during the same time may indicate that CSM interaction may be the reason for this discrepancy.
For our 56Ni mass estimates we use the prescriptions described in Hamuy (2003) and Jerkstrand et al. (2012) both of which employ the radioactive tail of the bolometric lightcurve. We estimate the value at various epochs and find that on day 103 we get 56Ni masses of 0.058 0.002 M☉ and 0.055 M☉ 0.002 respectively. These values fall to 56Ni masses of 0.050 0.003 M☉ and 0.047 M☉ 0.003 on day 153. We, therefore, estimate the 56Ni mass to be 0.05 M☉ with the caveat that extra heating from late CSM interaction can inflate this value and that incomplete gamma ray trapping is not taken into account.
We also calculate the temperature and photospheric radius using the Light Curve Fitting Package shown in the top panel of Figure 6. The first few epochs reveal a rise in temperature TBB from 15.0 kK on 1.7 days to a max of 23.4 kK on day 3.5. This is significantly cooler than the maximum temperatures of SN 2023ixf of 34.4 kK (Zimmerman et al., 2024) and 30.7 kK for SN 2024ggi (Shrestha et al., 2024a) but consistent with the 25 kK max for SN 2022jox (Andrews et al., 2024). Note that the cadence of Swift observations were different for all four SNe, which could contribute to some of the differences as the bulk of the luminosity is in the UV at this point. There is then a general decline of TBB from max over the next 60–80 days until reaching a value of 4.75 kK as it moves to the radioactive tail. The photospheric radius, RBB, rises to a maximum of roughly 2.9 104 R☉ during the first 60 or so days as the SN is on the plateau.
@@ 5.1 Optical Spectra
As summarized in Table 2 and shown in Figures 7 and 8, we obtained optical spectra from 2.9 days after explosion until day 147 when SN 2024bch was lost behind the Sun. On day 2.9 we have both low- and medium-resolution data, which show H I emission, along with the high-ionization lines of He II 4686 and C IV 5801,5811. Note in our higher-resolution GMOS B1200 data we are able to easily resolve the separate C IV lines as well as He II 6560 which is often blended with H. Absent are the lines of N III 4634,4641, C III 4648,4650 and N V 7109,7123 which has been seen in the early spectra of other Type II SNe. This does not discount their presence at earlier epochs, but by 2.9 days if they had existed they have faded away. In the day 1.40 and 1.42 spectra presented in Tartaglia et al. (2024) there is a feature blueward of He II 4686 that has been identified as N III; this is likely correct but the resolution makes it difficult to definitively confirm. These high ionization lines all disappear after our 3.9 day spectrum and before our 6 day spectrum. The Balmer emission lines persist for a few more days after the other high ionization lines are gone. Combining an ejecta velocity of 7000 km s-1 measured from H and Fe II 5169 from our day 16 and 22 spectra and the range of dates when the disappearance of high ionization lines occurred we can estimate an approximate CSM radius of RCSM = 2.4 – 3.6 1014 cm. Assuming a progenitor wind velocity of 35 km s-1 that we measure from the high resolution spectrum (and discussed in detail below), this quiescent mass loss would have occurred 2.4–3.6 years prior (much earlier than the 9 years derived in Tartaglia et al. (2024)). It is interesting to note though that no obvious outburst was seen during this time period in the ATLAS and ZTF precursor imaging.

### v2

@@ Abstract
We present a comprehensive multi-epoch photometric and spectroscopic study of SN 2024bch, a nearby (19.9 Mpc) Type II supernova (SN) with prominent early high ionization emission lines. Optical spectra from days after the estimated explosion reveal narrow lines of H I, He II, C IV, and N IV that disappear by day 6. High cadence photometry from the ground and TESS show that the SN brightened quickly and reached a peak M 17.8 mag within a week of explosion, and late-time photometry suggests a 56Ni mass of 0.050 M☉. High-resolution spectra from day and 43 trace the unshocked circumstellar medium (CSM) and indicate a wind velocity of 30–40 km s-1, a value consistent with a red supergiant (RSG) progenitor. Comparisons between models and the early spectra suggest a pre-SN mass-loss rate of , which is too high to be explained by quiescent mass loss from RSGs, but is consistent with some recent measurements of similar SNe. Persistent blueshifted H I and [O I] emission lines seen in the optical and NIR spectra could be produced by asymmetries in the SN ejecta, while the multi-component H may indicate continued interaction with an asymmetric CSM well into the nebular phase. SN 2024bch provides another clue to the complex environments and mass-loss histories around massive stars.
@@ 1 Introduction
Extensive mass-loss is an important property of massive stars, and in many cases it appears to ramp up to extreme levels in the years to decades before core collapse (see Smith, 2014, and references therein). While the final stages of a massive star’s life may be hard to observe, some of the information is recorded in the surrounding stellar envelope and circumstellar material (CSM) and can be observed during the first few hours and days of a core collapse supernova (CCSN) explosion. This ‘flash spectroscopy’ includes fleeting narrow lines of H, He, C, and N. While this phenomenon had been well documented in several lucky cases described in the literature (Niemela et al., 1985; Benetti et al., 1994; Leonard et al., 2000; Quimby et al., 2007; Shivvers et al., 2015), within the last decade the numbers of CCSN observed early enough to detect interaction with the mass-loss from the progenitor has dramatically increased thanks to fast response by dedicated transient surveys (for example: Gal-Yam et al., 2014; Smith et al., 2015; Yaron et al., 2017; Khazov et al., 2016; Bullivant et al., 2018; Hosseinzadeh et al., 2018; Tartaglia et al., 2021; Bruch et al., 2021; Terreran et al., 2022; Jacobson-Galán et al., 2022; Bruch et al., 2023; Jacobson-Galán et al., 2024a; Andrews et al., 2024). The recent and very nearby ( 7 Mpc) SNe 2023ixf and 2024ggi alone have contributed many publications analyzing their early-time luminosity evolution and exquisite early time spectroscopic datasets (Smith et al., 2023; Li et al., 2024; Zimmerman et al., 2024; Jacobson-Galán et al., 2023; Bostroem et al., 2023a; Hiramatsu et al., 2023; Teja et al., 2023; Zhang et al., 2023; Shrestha et al., 2024a; Jacobson-Galán et al., 2024b; Zhang et al., 2024; Pessi et al., 2024; Dickinson et al., 2024).
@@
The CCSNe that result from a progenitor that retains most of the hydrogen envelope prior to explosion are classified as Type II (see Arcavi, 2017; Gal-Yam, 2017; Branch & Wheeler, 2017, for detailed reviews). Some of these objects, particularly those with plateau light curves (SNe II-P), have been confirmed to come from red supergiant (RSG) progenitor stars via (Smartt, 2015; Van Dyk, 2015). When Type II SNe are observed early enough (hours to a few days after explosion) we can often observe shock interaction between the ejecta and the CSM or from the shock breakout of the SN itself (Bruch et al., 2021). These fleeting emission lines show Lorentzian profiles that can have wings extending out to 1500 - 2000 km s-1 due to electron scattering in the ionized gas (see Smith, 2017 for a review of SNe with dense CSM). Early spectroscopy can be a powerful device for understanding this enhanced mass loss and its potential mechanisms prior to collapse by probing the density, temperature, and composition of the surrounding CSM.
@@
While eruptive mass loss in luminous blue variables (LBVs) and strong line-driven winds of Wolf-Rayet (WR) stars are well known (see Smith, 2014 for a review), for RSGs there have been significant reductions in wind mass-loss prescriptions, and significant mismatch between these expected RSG wind mass-loss rates and mass-loss rates calculated from CSM interaction in SNe. For instance, the recent RSG wind mass-loss prescription from Beasor et al. (2020), which is estimated from modern high-quality IR data, reduces RSG mass-loss rates by about an order of magnitude compared to the older 1970s and 1980s literature-based canonical prescription from de Jager et al. (1988). Some stellar evolution models invoke even higher “enhanced” mass-loss rates (Ekström et al., 2012), but even these artificially inflated rates are still not enough to explain the estimated via flash spectroscopy observations (e.g., Jacobson-Galán et al., 2024a). This discrepancy could be resolved if periods of enhanced pre-SN mass loss are considered for RSGs rather than steady, quiescent mass loss from winds.
@@
Causes for increased periods of mass loss could be from wave driven mass loss during late-stage nuclear burning (Quataert & Shiode, 2012; Fuller, 2017; Wu & Fuller, 2021), turbulent convection in the core due to some dynamical instability (Smith & Arnett, 2014), or even interaction with a companion and a SN progenitor’s inflated envelope (Smith & Arnett, 2014; Soker, 2013), as the binary fraction for massive stars is high (Sana et al., 2012; Moe & Di Stefano, 2017). Eruptive mass loss from the progenitors of normal Type II SNe have been very difficult to detect observationally (Kochanek et al., 2017; Dong et al., 2023; Shrestha et al., 2024a, for example) suggesting if they occur they are likely some combination of dusty, faint, and quick, although see Jacobson-Galán et al. (2022) for one possible exception.
@@ 2 Discovery, Distance, and Reddening
SN 2024bch is located in NGC 3206 (Figure 1) at J2000 coordinates , . It was discovered on 2024-01-29.25 (Wiggins, 2024, MJD 60338.25) and classified later that night as a possible IIn (n is for “narrow” spectral lines), resembling the well-studied IIn SN 1998S (Balcon, 2024). While there are similarities between the two spectra, the further evolution of SN 2024bch indicated that its IIn-like features faded more quickly than SN 1998S, and that the narrow lines were likely those typically seen in objects exhibiting flash ionization. A nondetection of the object was reported by the Gravitational-wave Optical Transient Observer (Steeghs et al., 2022, GOTO) on 2024-01-28.04 (MJD 60337.04) with a limiting magnitude of 18.2 in their L-GOTO filter. A day later on 2024-01-29.06 (MJD 60338.06) GOTO recorded a detection at the same magnitude of L = 18.2 0.06, which was observed a few hours before the data cited in the discovery report. Without any other observations during this time to help constrain the date of explosion we will use the midway point between the GOTO limiting magnitude and the , to be our explosion epoch which we will use throughout this paper.
@@
| Discovery (JD) | 2460338.75 |
| Explosion Epoch (JD)aamid point of last nondetection and | |
@@ 4.1 Optical and UV Lightcurves
The full optical and UV lightcurves of SN 2024bch are shown in Figure 3, and the V-band lightcurve in absolute magnitude is shown in the bottom of Figure 4. Our estimated explosion epoch of MJD = is indicated by the dashed line in the right panel of Figure 3. We were able to observe fairly continuously from 1.5 days out to 155 days before the SN became Sun constrained. From our data it appears that the V band reaches a maximum mag at roughly 7 days post explosion, while the and -band takes much longer, roughly 15 to 17 days to reach the maximum. This could potentially be due to CSM interaction that will be discussed in detail below, but is still within somewhat normal values for some Type II SNe (Anderson et al., 2014b; González-Gaitán et al., 2015). The observations also show a peak at roughly 5 days, once again on the shorter end of the distribution, but within the normal range. We also catch a rise in the Swift data between our first and second epoch indicating maximum was reached somewhere near day 3. Unfortunately, the cadence is not very high for our Swift observations and our first epoch was at days after explosion so we cannot constrain the UV behavior more than this.
As can be seen in Figure 3, the plateau length persists only until 70 days post-explosion, which is a bit on the shorter side for Type II-P supernovae (Anderson et al., 2014b; Valenti et al., 2016). The fall from the plateau lasts about 15-20 days (the observational cadence was quite low during this time), and when the radioactive tail phase begins mag, or about 3 magnitudes fainter than at peak (bottom panel Figure 4).
Overall the shape and brightness of the V-band lightcurve is quite similar to the recent and well-studied Type II SN 2023ixf, both in plateau duration and shape and in the post-plateau brightness (Hsu et al., 2024). The lightcurves of SN 2024ggi and SN 2022jox, other recent CCSNe showing flash features, also have very similar properties at least until about 50 days when the lightcurve of SN 2024ggi shows a much more pronounced and prolonged plateau. SN 2017gmr and SN 2017eaw, and the second half of the plateau phase for SN 2024ggi, show the more classical plateau shape and length. Interestingly, both SN 2017gmr and SN 2017eaw were discovered quite early and have spectroscopy within the first few days and do not show flash features other than possibly fleeting narrow H and broad He bumps (Andrews et al., 2019; Rui et al., 2019), and SN 2024ggi only showed flash features until day 3-4 (Shrestha et al., 2024a; Zhang et al., 2024; Jacobson-Galán et al., 2024b).
@@ 4.2 Color Evolution
In the top panel of Figure 4 we show the color evolution of SN 2024bch compared with other well studied Type II SNe. . As seen from the figures, we are able to catch a blueward color evolution over the first three days, reaching mag, and then a slow evolution to the red until reaching a value of = 1.0 mag around day 75, when the fall from plateau begins. From this point onward there is a quick drop to = 0.9 mag and the color slowly trends blueward with a = 0.8 mag around day 150 when the SN becomes Sun constrained.
@@ 4.3 Bolometric Lightcurve and 56Ni Mass
The late-time bolometric luminosity appears to decline at a rate of 0.014 mag d-1, a value higher than the canonical rate of fully-trapped 56Co decay of 0.0098 mag d-1. This is shown as the dashed line in the bottom of Figure 6 where we show the decay of a mass of 56Ni = 0.055 M☉ (a value we derive below) for comparison. The slight deviation and quicker decline of SN 2024bch may indicate that the gamma rays were not fully trapped (Anderson et al., 2014b; Martinez et al., 2022), that there is potentially dust formation occurring, or that there is CSM interaction which is contributing to the overall luminosity and not only increasing the luminosity but fading at a higher rate than radioactive decay alone. While we will explore these in more detail later, average values for 56Ni masses in normal Type II SNe have been found to be 0.04 M☉ (Valenti et al., 2016; Müller et al., 2017; Anderson, 2019; Rodríguez et al., 2021). The slightly larger estimated 56Ni mass combined with the shape of the H emission (described in section 5.3) during the same time may indicate that CSM interaction may be the reason for this discrepancy.
For our 56Ni mass estimates we use the prescriptions described in Hamuy (2003) and Jerkstrand et al. (2012) both of which employ the radioactive tail of the bolometric lightcurve. We estimate the value at various epochs and find that on day 103 we get 56Ni masses of 0.058 M☉ 0.002 and 0.055 M☉ 0.002 respectively. These values fall to 56Ni masses of 0.050 M☉ 0.003 and 0.047 M☉ 0.003 on day 153. We, therefore, estimate the 56Ni mass to be 0.05 M☉ with the caveat that extra heating from late CSM interaction can inflate this value and that incomplete gamma ray trapping is not taken into account.
We also calculate the temperature and photospheric radius using the Light Curve Fitting Package shown in the top panel of Figure 6. The first few epochs reveal a rise in temperature TBB from 15.0 kK on days to a max of 23.4 kK on day . , this is significantly cooler than the maximum temperatures of SN 2023ixf of 34.4 kK (Zimmerman et al., 2024) and 30.7 kK for SN 2024ggi (Shrestha et al., 2024a) but consistent with the 25 kK max for SN 2022jox (Andrews et al., 2024). Note that the cadence of Swift observations were different for all four SNe, which could contribute to some of the differences as the bulk of the luminosity is in the UV at this point. There is then a general decline of TBB from max over the next 60–80 days until reaching a value of 4.75 kK as it moves to the radioactive tail. The photospheric radius, RBB, rises to a maximum of roughly 2.9 104 R☉ during the first 60 or so days as the SN is on the plateau.
@@ 5.1 Optical Spectra
As summarized in Table 2 and shown in Figures 7 and 8, we obtained optical spectra from days after explosion until day 147 when SN 2024bch was lost behind the Sun. On day we have both low- and medium-resolution data, which show H I emission, along with the high-ionization lines of He II 4686 and C IV 5801,5811. Note in our higher-resolution GMOS B1200 data we are able to easily resolve the separate C IV lines as well as He II 6560 which is often blended with H. Absent are the lines of N III 4634,4641, C III 4648,4650 and N V 7109,7123 which has been seen in the early spectra of other Type II SNe. This does not discount their presence at earlier epochs, but by days if they had existed they have faded away. In the day 1.40 and 1.42 spectra presented in Tartaglia et al. (2024) there is a feature blueward of He II 4686 that has been identified as N III; this is likely correct but the resolution makes it difficult to definitively confirm. These high ionization lines all disappear after our day spectrum and before our day spectrum. The Balmer emission lines persist for a few more days after the other high ionization lines are gone. Combining an ejecta velocity of 7000 km s-1 measured from H and Fe II 5169 from our day 16 and 22 spectra and the range of dates when the disappearance of high ionization lines occurred we can estimate an approximate CSM radius of RCSM = 2.4 – 3.6 1014 cm. Assuming a progenitor wind velocity of 35 km s-1 that we measure from the high resolution spectrum (and discussed in detail below), this quiescent mass loss would have occurred 2.4–3.6 years prior (much earlier than the 9 years derived in Tartaglia et al. (2024)). It is interesting to note though that no obvious outburst was seen during this time period in the ATLAS and ZTF precursor imaging.